OpenStudy (anonymous):

can some help me show me how to start this problem. If a flea can jump straight up to a height of 0.440 m, what is the initial speed as it leaves the ground?

OpenStudy (saifoo.khan):

v^2 = u^2 + 2as

OpenStudy (saifoo.khan):

v^2 = u^2 + 2as 0^2 = u^2 + 2(-9.81)(0.440)

OpenStudy (saifoo.khan):

@satellite73 : What mistake am i doing?

OpenStudy (saifoo.khan):

No, it's correct i guess..